# Adapt and Diffuse: Sample-adaptive Reconstruction via Latent Diffusion Models

### v2

@@ 3.2 Sample-adaptive Inference
Diffusion-based inverse problem solvers synthesize missing data that has been destroyed by the degradation process through diffusion. As shown in Figure 2, depending on the amount of missing information (easy vs. hard samples), the optimal number of diffusion steps may greatly vary. Too few steps may not allow the diffusion process to generate realistic details on the image, leaving the reconstruction overly smooth. On the other hand, diffusion-based solvers are known to hallucinate details that may be inconsistent with the ground truth signal, or even become unstable when too many reverse diffusion steps are applied. Authors in [7] observe that there always exists an optimal spot between 00 and $N$ diffusion steps that achieves the best reconstruction performance. We aim to automatically find this "sweet spot" on a sample-by-sample basis.
[Uncaptioned image] Figure 2: The optimal number of reverse diffusion steps varies depending on the severity of degradations. Fixing the number of steps results in over-diffusing some samples, whereas others could benefit from more iterations.
[Uncaptioned image] Figure 3: Effect of degradation on predicted severities: given a ground truth image corrupted by varying amount of blur, $\hat{\sigma}$ is a non-decreasing function of the blur amount.
@@
where $c\geq 0$ is a tuning parameter.
@@ 4 Experiments
Dataset – We perform experiments on CelebA-HQ ($256\times 256$) [19] and LSUN Bedrooms [50]. We match the training and validation splits used in [33], and set aside $200$ images from the validation split for testing. For comparisons involving the CelebA dataset and image domain score models we test on FFHQ [20], as pre-trained image-domain score models have been trained on the entire CelebA-HQ dataset, unlike LDMs.
@@
Degradations – We consider three degradations of diverging characteristics. Gaussian blur: we apply Gaussian blur with kernel size of $61$. In order to investigate the effect of strongly varying degradation severity on sampling techniques, we analyze two settings: (1) varying Gaussian blur, where we sample the kernel standard deviation uniformly on $[0,3]$ (00 corresponds to no blurring) and (2) fixed Gaussian blur where the kernel standard deviation is set to $3$. Nonlinear blur: we deploy GOPRO motion blur simulated by a neural network model from [47]. This is a nonlinear forward model due to the camera response function. We randomly sample nonlinear blur kernels for each image. Varying random inpainting: we randomly mask $70-80\%$ of the pixels, where the masking ratio is drawn uniformly. In all experiments, we add Gaussian noise to images in the $[0,1]$ range with noise std of $0.05$.
@@
Baseline solvers and comparison methods – Flash-Diffusion acts as a wrapper around any latent diffusion solver adding sample-adaptivity and acceleration to the baseline inverse problem solver. We experiment with the following baseline solvers: Latent-DPS [39], a straightforward adaptation of DPS [5] to the latent domain; GML-DPS and PSLD (linear-only) [34], algorithms that are extensions of Latent-DPS with improved performance; and ReSample [39], a state-of-the-art latent diffusion solver leveraging hard data consistency. We compare the baseline solvers to their adaptive version obtained from our framework, which we denote by Flash(baseline solver). Furthermore, we compare the results with DPS [5], a well-established diffusion solver that operates in the pixel space. As our framework requires fine-tuning to the specific degradation, we additionally compare with SwinIR [27], a strong Transformer-based supervised image restoration model trained on corrupted-clean data pairs. Finally, we show results of decoding the severity encoder’s $\hat{\bm{z}}$ estimate directly without diffusion, denoted by AE (autoencoded). Further details on comparison methods are in Appendix A.
@@
Models – We use pre-trained DMs from [11] for DPS and from [33] for LDMs [^1] https://github.com/CompVis/latent-diffusion. We fine-tune severity encoders from pre-trained LDM encoders and utilize a single convolution layer on top of $\hat{\bm{z}}$ to predict $\hat{\sigma}$. For more details on the experimental setup and hyperparameters, see Appendix A.
@@
Figure 4: Blur amount ($t$) vs. predicted degradation severity ($\hat{\sigma}$). Outliers indicate that the predicted degradation severity is not solely determined by the amount of blur. The bottom image is ’surprisingly easy’ to reconstruct, as it is overwhelmingly smooth with features close to those seen in the training set. The top image is ’surprisingly hard’, due to more high-frequency details and unusual features not seen during training. Points in red suggest that a given degradation severity may result from a wide range of blur levels (more details in Appendix C).
@@ 4.1 Severity encoding
In this section, we investigate properties of the predicted degradation severity $\hat{\sigma}$. We perform experiments on a $1k$-image subset of the validation split. First, we isolate the effect of degradation on $\hat{\sigma}$ (Fig. 3). We fix the clean image and apply increasing amount of Gaussian blur. We observe that $\hat{\sigma}$ is an increasing function of the blur amount applied to the image: heavier degradations on a given image result in higher predicted degradation severity. This implies that the severity encoder learns to capture the amount of information loss caused by the degradation.
Next, we investigate the relation between $\hat{\sigma}$ and the underlying degradation severity (Fig. 4, top). We parameterize the corruption level by $t$, where $t=0$ corresponds to no blur and $t=1$ corresponds to the highest blur level where in both cases we add Gaussian noise ($\sigma=0.05$). We vary the blur kernel width linearly for $t\in(0,1)$. We observe that the predicted $\hat{\sigma}$ severities strongly correlate with the corruption level. However, the existence of outliers suggest that factors other than the corruption level may also contribute to the predicted severities. The bottom image is predicted to be ’surprisingly easy’, as other images of the same corruption level are typically assigned higher predicted severities. This sample is overwhelmingly smooth, with a lack of fine details and textures, such as hair, present in other images. Moreover, the image shares common features with others in the training set. On the other hand, the top image is considered ’surprisingly difficult’, as it contains unexpected features and high-frequency details that are uncommon in the dataset. This example highlights the potential application of our technique to hard example mining and dataset distillation. More discussion on the different contributors to degradation severity is included in Appendix C, and the limitations of severity encoding is studied in Appendix D.
@@ 4.2 Sample-adaptive reconstruction
Comparison of latent diffusion solvers– Table 1 summarizes our experimental results comparing various latent diffusion solvers with and without our Flash-Diffusion wrapper. Reconstructed samples are depicted in Figure 5. First, we observe that adaptive solvers consistently outperform their baselines across all degradations, datasets and image quality metrics. The performance improvement is especially significant in experiments with highly varying degradation severity, such as varying Gaussian blur, nonlinear blur with randomly sampled kernel and random inpainting, further highlighting the importance of sample-adaptivity. Interestingly, we observe that even though ReSample achieves the best performance among baseline solvers by a large margin, the performance gap diminishes between adaptive solvers with Flash(Latent-DPS) often outperforming more advanced techniques. We hypothesize that the differences between baseline solvers stem from differences in the diffusion dynamics in the early and middle stages (chaotic and semantic stages in [39]) of sampling, a regime typically bypassed by Flash-Diffusion. In terms of sampling speedup, we obtain 8×10×8$\times$-10$\times$8 × - 10 × acceleration in average number of reverse diffusion steps across the test dataset compared to the non-adaptive baseline, without any drop in reconstruction quality. Furthermore, we observe that simply leveraging DDIM acceleration in the baseline solvers, given the same compute budget in inference time, cannot compete with the reconstruction quality obtained from Flash-Diffusion (details in Appendix E).
Comparison with other methods – In Table 1 (top), we observe that SwinIR, a state-of-the-art supervised restoration model achieves higher PSNR and SSIM than diffusion methods, however it falls short in terms of perceptual image quality. This phenomenon is due to the perception-distortion trade-off [3]: improving perceptual image quality is fundamentally at odds with distortion metrics. Furthermore, we highlight that our initial reconstructions obtained from the severity encoder are poorer quality than SwinIR reconstructions (compare AE vs. SwinIR), however diffusion samplers greatly improve over the perceptual quality of the initial reconstruction, eventually surpassing both SwinIR and all non-adaptive diffusion solver baselines.
@@
Importance of sample-adaptivity – To highlight the importance of sample-adaptivity, we compare our adaptive technique to a simple non-adaptive modification. Here, instead of predicting the starting time via severity encoding on a sample-by-sample basis, we set the starting time for all samples to $N^{\prime}$. For the sake of simplicity, we choose Flash(Latent-DPS) as our adaptive method, and vary the fixed start time between $10-500$ for the non-adaptive variant. We leverage the SwinIR reconstructions as initialization for the non-adaptive variant. We perform experiments on CelebA-HQ with varying Gaussian blur, the results are depicted in Fig. 6. We observe that our adaptive method achieves the best FID across any fixed number of steps by a large margin, despite the lower quality initialization from the severity encoder. Moreover, our technique with sample-adaptivity achieves near-optimal LPIPS with 2×2$\times$2 × less average number of diffusion steps compared to the non-adaptive variant. We observe that the predicted start times from our technique are spread around the mean and not closely concentrated, further highlighting the adaptivity of Flash-Diffusion.
Robustness and limitations– The performance of Flash-Diffusion relies on the accuracy of the severity encoder trained in a supervised fashion on the degraded image distribution. Therefore, it is crucial to understand the robustness of severity encoding to noise level and forward model shifts in test time. In terms of noise robustness, we observe that severity encoding maintains its performance on noise levels lower than in the training setting, however it breaks down at noise levels well above the training regime. In terms of forward model mismatch, we find that as long as the train and test forward operators are similar (heavy Gaussian blur and nonlinear blur), severity encoding can maintain acceptable performance. However in case of significant mismatch, the accuracy of severity encoding breaks down. An in-depth characterization of severity encoder robustness is detailed in Appendix D.
Moreover, we analyze the robustness of Flash-Diffusion solvers compared to their non-adaptive counterparts with respect to a mismatch in forward model noise level. In particular, we base our comparison on ReSample, the best performing baseline solver in our experiments. Fig. 7 demonstrates that the performance of ReSample degrades more rapidly than its adaptive counterpart when the noise level is higher than expected, following a similar trend to SwinIR, a supervised non-diffusion reconstructor. In the low-noise regime, Flash(ReSample) shows improved robustness compared to the baseline solver. More experiments on forward model mismatch can be found in Appendix D.2.
@@
Figure 6: Comparison of Flash(Latent-DPS) with the non-adaptive variant with fixed starting time for all samples (varying Gaussian blur, CelebA-HQ). Left and center: We plot reconstruction quality as a function of the number of diffusion steps for the non-adaptive variant with various fixed starting times. The data point corresponding to our adaptive method depicts the average number of diffusion steps due to adaptive starting time. Our adaptive method achieves the best FID and near-optimal LPIPS compared to any choice of non-adaptive starting time. Right: Histogram of predicted starting times for Flash(Latent-DPS).
@@ 5 Conclusions
In this work, we make the key observation that the difficulty of solving an inverse problem may vary greatly on a sample-by-sample basis, depending on the ground truth signal structure, the applied corruption, the model, the training set and the complex interactions between these factors. Despite this natural variation in the difficulty of a reconstruction task, most techniques apply a rigid model that expends the same compute irrespective of the amount of information destroyed in the noisy measurements, resulting in suboptimal performance and wasteful resource allocation. We propose Flash-Diffusion, a sample-adaptive method that predicts the degradation severity of corrupted signals, and utilizes this estimate to automatically tune the compute allocated to reconstruct the sample. In particular, we use the prediction error of an encoder in latent space as a proxy for reconstruction difficulty, which we call severity encoding. Then, we leverage a latent diffusion process to reconstruct the sample, where the length of the sampling trajectory is directly scaled by the predicted severity. Our method can be coupled with any latent diffusion solver, greatly enhancing its reconstruction performance and accelerating sampling by up to 10×10$\times$10 ×. We demonstrate the gains in terms of reconstruction quality and speedup on a variety of degradations and datasets.

### v3

@@ 3.2 Sample-adaptive inference
Diffusion-based inverse problem solvers synthesize missing data that has been destroyed by the degradation process through diffusion. As shown in Figure 3, depending on the amount of missing information (easy vs. hard samples), the optimal number of diffusion steps may greatly vary. Too few steps may not allow the diffusion process to generate realistic details on the image, leaving the reconstruction overly smooth. On the other hand, diffusion-based solvers are known to hallucinate details that may be inconsistent with the ground truth signal, or even become unstable when too many reverse diffusion steps are applied. Authors in Chung et al. (2022c) observe that there always exists an optimal spot between 00 and $N$ diffusion steps that achieves the best reconstruction performance. We aim to automatically find this "sweet spot" on a sample-by-sample basis.
Figure 2: The optimal number of reverse diffusion steps varies depending on the severity of degradations. Fixing the number of steps results in over-diffusing some samples, whereas others could benefit from more iterations.
Figure 3: Effect of degradation on predicted severities: given a ground truth image corrupted by varying amount of blur, $\hat{\sigma}$ is a non-decreasing function of the blur amount.
@@
where $c\geq 0$ is a tuning parameter. We discuss alternatives to this initialization scheme based on adaptive DDIM encoding in Appendix E.
@@ 4 Experiments
Figure 4: Blur amount ($t$) vs. predicted degradation severity ($\hat{\sigma}$). Outliers indicate that the predicted degradation severity is not solely determined by the amount of blur. The bottom image is surprisingly easy to reconstruct, as it is overwhelmingly smooth with features close to those seen in the training set. The top image is surprisingly hard, due to more high-frequency details and unusual features not seen during training. Points in red suggest that a given degradation severity may result from a wide range of blur levels (see Fig. 5 and discussion under Identifying contributors to severity in Section 4.1).
@@
Dataset – We perform experiments on CelebA-HQ ($256\times 256$) (Karras et al., 2018) and LSUN Bedrooms (Yu et al., 2015). We match the training and validation splits used in Rombach et al. (2022), and set aside $200$ images from the validation split for testing. For comparisons involving the CelebA dataset and image domain score models we test on FFHQ (Karras et al., 2019), as pre-trained image-domain score models have been trained on the entire CelebA-HQ.
@@
Degradations – We consider three degradations of diverging characteristics. Gaussian blur: we apply Gaussian blur with kernel size of $61$. In order to investigate the effect of strongly varying degradation severity on sampling techniques, we analyze two settings: (1) varying Gaussian blur, where we sample the kernel standard deviation uniformly on $[0,3]$ (00 corresponds to no blurring) and (2) fixed Gaussian blur where the kernel standard deviation is set to $3$. Nonlinear blur: we deploy GOPRO motion blur simulated by a neural network model from Tran et al. (2021). This is a nonlinear forward model due to the camera response function. We randomly sample nonlinear blur kernels for each image. Varying random inpainting: we randomly mask $70-80\%$ of the pixels, where the masking ratio is drawn uniformly. In all experiments, we add Gaussian noise to images in the $[0,1]$ range with noise std of $0.05$.
@@
Baseline solvers and comparison methods – Flash-Diffusion acts as a wrapper around any latent diffusion solver adding sample-adaptivity and acceleration to the baseline inverse problem solver. We experiment with the following baseline solvers: Latent-DPS (Rout et al., 2024), a straightforward adaptation of DPS (Chung et al., 2022a) to the latent domain; GML-DPS and PSLD (linear-only) (Rout et al., 2024), algorithms that are extensions of Latent-DPS with improved performance; and ReSample (Song et al., 2023), a state-of-the-art latent diffusion solver leveraging hard data consistency. We compare the baseline solvers to their adaptive version obtained from our framework, which we denote by Flash(baseline solver). Furthermore, we compare the results with DPS (Chung et al., 2022a), a well-established diffusion solver that operates in the pixel space. As our framework requires fine-tuning to the specific degradation, we additionally compare with SwinIR (Liang et al., 2021), a strong Transformer-based supervised image restoration model trained on corrupted-clean data pairs. Finally, we show results of decoding the severity encoder’s $\hat{\bm{z}}$ estimate directly without diffusion, denoted by AE (autoencoded). Further details on comparison methods are in Appendix A.
@@
Models – We use pre-trained DMs from Dhariwal & Nichol (2021) for DPS and from Rombach et al. (2022) for LDMs. We fine-tune severity encoders from pre-trained LDM encoders and utilize a single convolution layer on top of $\hat{\bm{z}}$ to predict $\hat{\sigma}$. For more details on the experimental setup and hyperparameters, see Appendix A.
@@ 4.1 Severity encoding
In this section, we investigate properties of the predicted degradation severity $\hat{\sigma}$. We perform experiments on a $1k$-image subset of the validation split.
Effect of degradation level – First, we isolate the effect of degradation on $\hat{\sigma}$ (Fig. 3). We fix the clean image and apply increasing amount of Gaussian blur. We observe that $\hat{\sigma}$ is an increasing function of the blur amount applied to the image: heavier degradations on a given image result in higher predicted degradation severity. This implies that the severity encoder learns to capture the amount of information loss caused by the degradation.
Interaction between degradation and image – Next, we investigate the relation between $\hat{\sigma}$ and the underlying degradation severity (Fig. 4). We parameterize the corruption level by $t$, where $t=0$ corresponds to no blur and $t=1$ corresponds to the highest blur level where in both cases we add Gaussian noise ($\sigma=0.05$). We vary the blur kernel width linearly for $t\in(0,1)$. We observe that the predicted $\hat{\sigma}$ severities strongly correlate with the corruption level. However, the existence of outliers suggest that factors other than the corruption level may also contribute to the predicted severities. The bottom image is predicted to be surprisingly easy, as other images of the same corruption level are typically assigned higher predicted severities. This sample is overwhelmingly smooth, with a lack of fine details and textures, such as hair, present in other images. Moreover, the image shares common features with others in the training set. On the other hand, the top image is considered surprisingly difficult, as it contains unexpected features and high-frequency details that are uncommon in the dataset. This example highlights the potential application of our technique to hard example mining and dataset distillation.
Identifying contributors to severity – Finally, we analyze the contribution of different factors to the predicted degradation severities (Fig. 5). We apply severity encoding to both the clean image with noise (no blur) and the noisy and degraded image, resulting in predicted severities $\hat{\sigma}_{noisy}$ and $\hat{\sigma}_{degr.+noisy}$. We quantify the difficulty of samples relative to each other via percentiles of the above two quantities, where we use $\hat{\sigma}_{noisy}$ as a proxy for the difficulty originating from the image structure. The reason for not using the clean image without additive noise is due to the fact that the severity encoder is trained on varying degradation severities, but with a fixed additive noise. Therefore, the model is not intended to accept noiseless measurements as inputs. We observe that for a fixed $\hat{\sigma}_{degr.+noisy}$, the composition of degradation severity may greatly vary. The two presented images have been assigned approximately the same $\hat{\sigma}_{degr.+noisy}$, however the top image is inherently easy to encode (low $\hat{\sigma}_{noisy}$ percentile) compared to other images in the dataset, therefore the severity is mostly attributed to the image degradation. On the other hand, the bottom image with the same $\hat{\sigma}_{degr.+noisy}$ is less corrupted by blur, but with high $\hat{\sigma}_{noisy}$ indicating a difficult image. This example further corroborates the interaction between ground truth signal structure and the applied corruption in determining the difficulty of a reconstruction task. For a detailed discussion on the limitations of severity encoding, we refer the reader to Appendix C.
Figure 5: Contributors to severity. Degraded images with approx. the same $\hat{\sigma}$ may have different factors contributing to the predicted severity. The main contributor to $\hat{\sigma}$ in the top image is the image degradation (blur), whereas the bottom image is inherently more difficult to reconstruct.
@@ 4.2 Sample-adaptive reconstruction
Comparison of latent diffusion solvers– Table 1 summarizes our experimental results comparing various latent diffusion solvers with and without our Flash-Diffusion wrapper. Reconstructed samples are depicted in Figure 6. First, we observe that adaptive solvers consistently outperform their baselines across all degradations, datasets and image quality metrics. The performance improvement is especially significant in experiments with highly varying degradation severity, such as varying Gaussian blur, nonlinear blur with randomly sampled kernel and random inpainting, further highlighting the importance of sample-adaptivity. Interestingly, we observe that even though ReSample achieves the best performance among baseline solvers by a large margin, the performance gap diminishes between adaptive solvers with Flash(Latent-DPS) often outperforming more advanced techniques. We hypothesize that the differences between baseline solvers stem from differences in the diffusion dynamics in the early and middle stages (chaotic and semantic stages in Song et al. (2023)) of sampling, a regime typically bypassed by Flash-Diffusion. In terms of sampling speedup, we obtain 8×10×8$\times$-10$\times$8 × - 10 × acceleration in average number of reverse diffusion steps across the test dataset compared to the non-adaptive baseline, without any drop in reconstruction quality. Furthermore, we observe that simply leveraging DDIM acceleration in the baseline solvers, given the same compute budget in inference time, cannot compete with the reconstruction quality obtained from Flash-Diffusion (details in Appendix D).
@@
Figure 7: Comparison of Flash(Latent-DPS) with a non-adaptive variant with fixed starting time (varying Gaussian blur, CelebA-HQ). Top: Reconstruction quality as a function of the number of diffusion steps for the non-adaptive variant with various fixed starting times. The data point corresponding to our adaptive method depicts the average number of diffusion steps due to adaptive starting time. Our adaptive method achieves the best FID and near-optimal LPIPS compared to any choice of non-adaptive starting time. Bottom: Histogram of predicted starting times for Flash(Latent-DPS).
Comparison with other methods – In Table 1 (top), we observe that SwinIR, a state-of-the-art supervised restoration model achieves higher PSNR and SSIM than diffusion methods, however it falls short in terms of perceptual image quality. This phenomenon is due to the perception-distortion trade-off (Blau & Michaeli, 2018): improving perceptual image quality is fundamentally at odds with distortion metrics. Furthermore, we highlight that our initial reconstructions obtained from the severity encoder are poorer quality than SwinIR reconstructions (compare AE vs. SwinIR), however diffusion samplers greatly improve over the perceptual quality of the initial reconstruction, eventually surpassing both SwinIR and all non-adaptive diffusion solver baselines.
Importance of sample-adaptivity – In order to highlight the importance of sample-adaptivity, we compare our adaptive technique to a simple non-adaptive modification. Here, instead of predicting the starting time via severity encoding on a sample-by-sample basis, we set the starting time for all samples to $N^{\prime}$. For the sake of simplicity, we choose Flash(Latent-DPS) as our adaptive method, and vary the fixed start time between $10-500$ for the non-adaptive variant. We leverage SwinIR reconstructions as initialization for the non-adaptive variant. We perform experiments on CelebA-HQ under varying Gaussian blur. The results are depicted in Fig. 7. Our adaptive method achieves the best FID across any fixed number of steps by a large margin, despite the lower quality initialization from the severity encoder. Moreover, our technique with sample-adaptivity achieves near-optimal LPIPS with 2×2$\times$2 × less average number of diffusion steps compared to the non-adaptive variant. We observe that the predicted start times from our technique are spread around the mean and not closely concentrated, further highlighting the adaptivity of Flash-Diffusion.
Robustness and limitations– The performance of Flash-Diffusion relies on the accuracy of the severity encoder trained in a supervised fashion on the degraded image distribution. Therefore, it is crucial to understand the robustness of severity encoding to noise level and forward model shifts in test time. In terms of noise, we observe that severity encoding maintains its performance on noise levels lower than in the training setting, however it breaks down at noise levels well above the training regime. In terms of forward model mismatch, we find that as long as the train and test operators are similar (heavy Gaussian blur and nonlinear blur), severity encoding can maintain acceptable performance. However in case of significant mismatch, the accuracy of severity encoding breaks down. An in-depth characterization of severity encoder robustness is detailed in Appendix C.
Moreover, we analyze the robustness of Flash-Diffusion solvers compared to their non-adaptive counterparts with respect to a mismatch in forward model noise level. In particular, we base our comparison on ReSample, the best performing baseline solver in our experiments. Fig. 8 demonstrates that the performance of ReSample degrades more rapidly than its adaptive counterpart when the noise level is higher than expected, following a similar trend to SwinIR, a supervised non-diffusion reconstructor. In the low-noise regime, Flash(ReSample) shows improved robustness compared to the baseline solver. More experiments on forward model mismatch can be found in Appendix C.2.
@@ 5 Conclusions
In this work, we make the key observation that the difficulty of solving an inverse problem may vary greatly on a sample-by-sample basis, depending on the ground truth signal structure, the applied corruption, the model, the training set and the complex interactions between these factors. Despite this natural variation in the difficulty of a reconstruction task, most techniques apply a rigid model that expends the same compute irrespective of the amount of information destroyed in the noisy measurements, resulting in suboptimal performance and wasteful resource allocation. We propose Flash-Diffusion, a sample-adaptive method that predicts the degradation severity of corrupted signals, and utilizes this estimate to automatically tune the compute allocated to reconstruct the sample. In particular, we use the prediction error of an encoder in latent space as a proxy for reconstruction difficulty, which we call severity encoding. Then, we leverage a latent diffusion process to reconstruct the sample, where the length of the sampling trajectory is directly scaled by the predicted severity. Our method can be coupled with any latent diffusion solver, greatly enhancing its performance and accelerating sampling by up to 10×10$\times$10 ×. We demonstrate the gains in terms of reconstruction quality and speedup on a variety of degradations and datasets.
Impact Statement
The presented paper leverages deep learning for reconstructing corrupted images. Deep learning-based reconstruction approaches are known to hallucinate realistic-looking features that may be inconsistent with the ground truth. Therefore, deploying our technique in safety-critical applications requires additional caution. Moreover, the technique leverages diffusion models capable of generating photorealistic synthetic images, which in turn can be abused by bad actors to spread misinformation.
Acknowledgments
We thank the anonymous reviewers who provided valuable suggestions and constructive feedback on our work. M. Soltanolkotabi is supported by the Packard Fellowship in Science and Engineering, a Sloan Research Fellowship in Mathematics, an NSF-CAREER under award #1846369, DARPA FastNICS program, and NSF-CIF awards #1813877 and #2008443. and NIH DP2LM014564-01. This research is also in part supported by AWS credits through an Amazon Faculty research award.